# Repetita Iuvant: Data Repetition Allows SGD to Learn High-Dimensional Multi-Index Functions

### v1

@@ A.3 Proof of Theorem 3
A.4 Hardness of sign functions
@@
$h^{\star}(\bm{x})=\mathrm{sign}(x_{1}\dots x_{m})$
@@
We first show the lower bound. For any vector $\bm{v}\in\mathbb{R}^{m}$, and $k<m$, $H_{k}(\langle\bm{v},\bm{x}\rangle)$ is a polynomial in $z$ of degree $k$, and therefore each monomial term is missing at least one variable. We show that this implies the lower bound of Proposition 3 through the following lemma:
@@
$\mathbb{E}\left[f(h^{\star}(\bm{x}))g(\bm{x})\right]=\frac{f(1)+f(-1)}{2}\mathbb{E}\left[g(\bm{x})\right]$
@@
$\mathbb{E}\left[f(h^{\star}(\bm{x}))g(\bm{x})\right]$ $=\mathbb{E}\left[\mathbb{E}\left[f(h^{\star}(\bm{x}))g(\bm{x})\mid x_{1},\dots,x_{m-1}\right]\right]$
$=\mathbb{E}\left[g(\bm{x})\mathbb{E}\left[f(h^{\star}(\bm{x}))\mid x_{1},\dots,x_{m-1}\right]\right]$
$=\mathbb{E}\left[g(\bm{x})\frac{f(1)+f(-1)}{2}\right]$
$=\frac{f(1)+f(-1)}{2}\mathbb{E}\left[g(\bm{x})\right],$
@@
Using this property on every monomial in $H_{k}(\langle\bm{v},\bm{x}\rangle)$ and summing back, we have for any function $f$
$\mathbb{E}\left[f(h^{\star}(x))H_{k}(\langle\bm{v},\bm{x}\rangle)\right]=\frac{f(1)+f(-1)}{2}\mathbb{E}\left[H_{k}(\langle\bm{v},\bm{x}\rangle)\right]$ (22)
But the latter is zero when $k\geq 1$ by orthogonality of the Hermite polynomials. This shows that $\ell_{\bm{v}}^{\star}\geq m$ for all $\bm{v}\in\mathbb{R}^{k}$, and therefore $m\leq\ell^{\star}\leq\ell$.
For the upper bound, let $\bm{v}=\bm{e}_{1}+\dots+\bm{e_{m}}$. Then
$H_{k}(\langle\bm{v},\bm{x}\rangle)=m!\,x_{1}\dots x_{m}+Q(\bm{x}),$
@@
$\mathbb{E}\left[h^{\star}(\bm{x})H_{k}(\langle\bm{v},\bm{x}\rangle)\right]$ $=m!\,\mathbb{E}\left[h^{\star}(\bm{x})x_{1}\dots x_{m}\right]$
@@
Appendix B Additional Numerical Investigation
@@
B.1 Testing wider range of Algorithm 1
@@
Figure 5: example of two different algorithm with data repetitions learning an hard single-index target $h^{\star}(\bm{z})=\mathrm{He}_{3}(z_{1})$, $\ell=3,\ell^{\star}=1$. Left: SAM, with $\rho_{0}=0.1,\gamma_{0}=0.01$. Right: 2-Lookahed with $\gamma_{0}=0.1$. See details in App. C.
@@
B.2 Extensive batch size
@@
Note that this drastically reduces the time steps needed to learn the teacher’s directions, although the sample complexity does not change. Dandi et al. [2024] first shows that some high-information exponent functions, such as $\mathrm{He}_{3}$, can be learned in just a few steps with full-batch gradient descent $n_{b}=O(d)$. In the same work, they introduce a larger class of hard functions that cannot be recovered in $T=O(1)$, not providing any information on possible bounds on the number of time steps, nor any claim suggesting that not all the function in the class have the same sample complexity. Our theory can provide a full understanding of the phenomena:
  • $\mathrm{He}_{3}$ has generative exponent $\ell^{\star}=1$, that implies we need $O(d)$ samples to weakly recover the target; when training with batch size $n_{b}=O(d)$, we have $T=O(1)$ as in Dandi et al. [2024]; Figure 6 (Left) is a numerical proof of this fact.
@@
Figure 6: Single-index recovery for large-batch training. The dashed horizontal line $\frac{1}{\sqrt{d}}$ is a visual guide to place random performance. The plots show $\sigma=\mathrm{relu}$, $\gamma=0.01$, $\rho=0.1$, and the performance is computed averaging across 10 different runs. Left $(\ell,\ell^{\star})=(3,1)$: $h^{\star}=\mathrm{He}_{3}.$ Right $(\ell,\ell^{\star})=(4,2)$: $h^{\star}=\mathrm{He}_{4}$. See details in App. C.
@@
Figure 7: Multi-index recovery for large-batch training. The dashed horizontal line $\frac{1}{\sqrt{d}}$ is a visual guide to place random performance. The plots show $\sigma=\operatorname{relu}$. Upper $(\ell_{\bm{v}},\ell_{\bm{v}}^{\star})=(3,1)$: $h^{\star}(\bm{z})=\operatorname{sign}(z_{1}z_{2}).$ Lower $(\ell_{\bm{v}},\ell_{\bm{v}}^{\star})=(4,2)$: $h^{\star}(\bm{z})=z_{1}z_{2}z_{3}$. See Appendix C for details.
Appendix C Implementation Details
@@
Unless explicitly stated, we always run the plain version of Algorithm 1 with squared loss, as opposed to Assumption 1 which requires correlation loss and spherical gradient. Also, the activation function we always use is ReLU, which does not satisfy Assumption 2, but is closer to real cases. Finally, we always run simulations with $W^{0}\bot W^{\star}$, which does not fall under Assumption 3. Indeed, our numerical simulations deviate from the theoretical assumptions needed for the formal proofs, but the aim here is to show that our claims are valid beyond the technical limitations of the theory and that they can allow us to understand settings closer to what is used in practice.
In the plots with a single-index target we plot the absolute value of the cosine similarity between the student weight $\bm{w}$ and the teacher weight $\bm{w}^{\star}$, averaged across $N$ different runs (where we vary both seen samples and initial conditions). If $q$ in the index for different runs:
$\text{yaxis}(t)=\frac{1}{N}\sum_{q=1}^{N}\left|\mathrm{CosSim}(\bm{w}^{(q)}_{t},\bm{w}^{\star{(q)}})\right|,$
@@
$\mathrm{CosSim}(\bm{w},\bm{w}^{\star})=\frac{\bm{w}\cdot\bm{w}^{\star}}{\left\lVert\bm{w}\right\rVert_{2}\left\lVert\bm{w}^{\star}\right\rVert_{2}}.$
@@
$\text{yaxis}(t)=\frac{1}{N}\sum_{q=1}^{N}\max_{j\in[p]}\left|\mathrm{CosSim}(\bm{w}^{(q)}_{j,t},\bm{w}^{\star{(q)}})\right|.$
C.1 Hyperparameters of the Figures

### v2

@@ C.2 Proof of Theorem 3
C.3 Hardness of sign functions
@@
$h^{\star}(\mathbf{x})=\mathrm{sign}(x_{1}\dots x_{m})$
@@
We first show the lower bound. For any vector $\mathbf{v}\in\mathbb{R}^{m}$, and $k<m$, $H_{k}(\langle\mathbf{v},\mathbf{x}\rangle)$ is a polynomial in $z$ of degree $k$, and therefore each monomial term is missing at least one variable. We show that this implies the lower bound of Proposition 6 through the following lemma:
@@
$\mathbb{E}\left[f(h^{\star}(\mathbf{x}))g(\mathbf{x})\right]=\frac{f(1)+f(-1)}{2}\mathbb{E}\left[g(\mathbf{x})\right]$
@@
$\mathbb{E}\left[f(h^{\star}(\mathbf{x}))g(\mathbf{x})\right]$ $=\mathbb{E}\left[\mathbb{E}\left[f(h^{\star}(\mathbf{x}))g(\mathbf{x})\mid x_{1},\dots,x_{m-1}\right]\right]$
$=\mathbb{E}\left[g(\mathbf{x})\mathbb{E}\left[f(h^{\star}(\mathbf{x}))\mid x_{1},\dots,x_{m-1}\right]\right]$
$=\mathbb{E}\left[g(\mathbf{x})\frac{f(1)+f(-1)}{2}\right]$
$=\frac{f(1)+f(-1)}{2}\mathbb{E}\left[g(\mathbf{x})\right],$
@@
Using this property on every monomial in $H_{k}(\langle\mathbf{v},\mathbf{x}\rangle)$ and summing back, we have for any function $f$
$\mathbb{E}\left[f(h^{\star}(x))H_{k}(\langle\mathbf{v},\mathbf{x}\rangle)\right]=\frac{f(1)+f(-1)}{2}\mathbb{E}\left[H_{k}(\langle\mathbf{v},\mathbf{x}\rangle)\right]$ (45)
But the latter is zero when $k\geq 1$ by orthogonality of the Hermite polynomials. This shows that $\ell_{\mathbf{v}}^{\star}\geq m$ for all $\mathbf{v}\in\mathbb{R}^{k}$, and therefore $m\leq\ell^{\star}\leq\ell$.
For the upper bound, let $\mathbf{v}=\mathbf{e}_{1}+\dots+\mathbf{e_{m}}$. Then
$H_{k}(\langle\mathbf{v},\mathbf{x}\rangle)=m!\,x_{1}\dots x_{m}+Q(\mathbf{x}),$
@@
$\mathbb{E}\left[h^{\star}(\mathbf{x})H_{k}(\langle\mathbf{v},\mathbf{x}\rangle)\right]$ $=m!\,\mathbb{E}\left[h^{\star}(\mathbf{x})x_{1}\dots x_{m}\right]$
@@
Appendix D Additional Numerical Investigation
@@
D.1 Testing wider range of Algorithm 1
@@
Figure 5: example of two different algorithm with data repetitions learning an hard single-index target $h^{\star}(\mathbf{z})=\mathrm{He}_{3}(z_{1})$, $\ell=3,\ell^{\star}=1$. Left: SAM, with $\rho_{0}=0.1,\gamma_{0}=0.01$. Right: 2-Lookahed with $\gamma_{0}=0.1$. See details in App. E.
@@
D.2 Extensive batch size
@@
Note that this drastically reduces the time steps needed to learn the teacher’s directions, although the sample complexity does not change. Dandi et al. [2024b] first shows that some high-information exponent functions, such as $\mathrm{He}_{3}$, can be learned in just a few steps with full-batch gradient descent $n_{b}=O(d)$. In the same work, they introduce a larger class of hard functions that cannot be recovered in $T=O(1)$, not providing any information on possible bounds on the number of time steps, nor any claim suggesting that not all the function in the class have the same sample complexity. Our theory can provide a full understanding of the phenomena:
  • $\mathrm{He}_{3}$ has generative exponent $\ell^{\star}=1$, that implies we need $O(d)$ samples to weakly recover the target; when training with batch size $n_{b}=O(d)$, we have $T=O(1)$ as in Dandi et al. [2024b]; Figure 6 (Left) is a numerical proof of this fact.
@@
Figure 6: Single-index recovery for large-batch training. The dashed horizontal line $\frac{1}{\sqrt{d}}$ is a visual guide to place random performance. The plots show $\sigma=\mathrm{relu}$, $\gamma=0.01$, $\rho=0.1$, and the performance is computed averaging across 10 different runs. Left $(\ell,\ell^{\star})=(3,1)$: $h^{\star}=\mathrm{He}_{3}.$ Right $(\ell,\ell^{\star})=(4,2)$: $h^{\star}=\mathrm{He}_{4}$. See details in App. E.
@@
Figure 7: Multi-index recovery for large-batch training. The dashed horizontal line $\frac{1}{\sqrt{d}}$ is a visual guide to place random performance. The plots show $\sigma=\operatorname{relu}$. Upper $(\ell_{\mathbf{v}},\ell_{\mathbf{v}}^{\star})=(3,1)$: $h^{\star}(\mathbf{z})=\operatorname{sign}(z_{1}z_{2}).$ Lower $(\ell_{\mathbf{v}},\ell_{\mathbf{v}}^{\star})=(4,2)$: $h^{\star}(\mathbf{z})=z_{1}z_{2}z_{3}$. See Appendix E for details.
Appendix E Implementation Details
@@
Unless explicitly stated, we always run the plain version of Algorithm 1 with squared loss, as opposed to Assumption 3 which requires correlation loss and spherical gradient. We use the reLU function as an activation. Finally, we always run simulations with $W^{0}\bot W^{\star}$, which does not fall under Assumption 2. Indeed, our numerical simulations deviate from the theoretical assumptions needed for the formal proofs, but the aim here is to show that our claims are valid beyond the technical limitations of the theory and that they can allow us to understand settings closer to what is used in practice.
In the plots with a single-index target we plot the absolute value of the cosine similarity between the student weight $\mathbf{w}$ and the teacher weight $\mathbf{w}^{\star}$, averaged across $N$ different runs (where we vary both seen samples and initial conditions). If $q$ in the index for different runs:
$\text{yaxis}(t)=\frac{1}{N}\sum_{q=1}^{N}\left|\mathrm{CosSim}(\mathbf{w}^{(q)}_{t},\mathbf{w}^{\star{(q)}})\right|,$
@@
$\mathrm{CosSim}(\mathbf{w},\mathbf{w}^{\star})=\frac{\mathbf{w}\cdot\mathbf{w}^{\star}}{\left\lVert\mathbf{w}\right\rVert_{2}\left\lVert\mathbf{w}^{\star}\right\rVert_{2}}.$
@@
$\text{yaxis}(t)=\frac{1}{N}\sum_{q=1}^{N}\max_{j\in[p]}\left|\mathrm{CosSim}(\mathbf{w}^{(q)}_{j,t},\mathbf{w}^{\star{(q)}})\right|.$
E.1 Hyperparameters of the Figures